OpenStudy (anonymous):

what is the y-coordinate of the solution of the system of equations? 3x+2y=-25 -x+y=0

OpenStudy (anonymous):

solve the second equation for x. Let me know what you get..

OpenStudy (anonymous):

idk....

OpenStudy (anonymous):

if you add something to each side of an equation does anything really change, besides the way the equation looks?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

7x=7 if x=1 then this equation is true, correct? if we add 3 to each side 7x+3=10 and x is still 1, the equation is still true just looks different

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

we want to get x by itself and positive the easiest way to do that is to add x to each side of the equation. y - x = 0 y - x + x = 0 + x

OpenStudy (anonymous):

y=x or x=y doesn't matter how you write it

OpenStudy (anonymous):

Now that we know this. we can substitute the x's for a y. This is called the substitution method. 3(y) + 2y = -25 Now solve for y